OpenStudy (anonymous):

Write a proof by mathmatical induction of the statement: 1^2+3^2+5^2+...+(2n-1)^2=(n(2n-1)(2n+1))/3

OpenStudy (anonymous):

anyone know?

OpenStudy (loser66):

your stuff is something, man!! no answer yet!!

OpenStudy (zarkon):

what have you tried

OpenStudy (anonymous):

i keep getting stuck. I really have no idea if im starting right or what.

OpenStudy (zarkon):

this is a fairly standard problem. what is the first thing you did?

OpenStudy (zarkon):

you get 10=10 with n=2

OpenStudy (loser66):

how" (2n-1)^2 = (2.2 -1)^2 = 3^2 =9

OpenStudy (zarkon):

\[1^2+3^2=1+9=10\]

OpenStudy (zarkon):

the LHS is a sum

OpenStudy (zarkon):

are you waiting for someone to do the problem for you?

OpenStudy (anonymous):

I would like to know how to go about solving the problem, or at least get a good link to an online guide.

OpenStudy (zarkon):

first do the basis step

OpenStudy (anonymous):

I have a solution and it has all steps. I just don't know how to let the student fill in some of the blanks. I'd have to give the whole thing.

OpenStudy (anonymous):

i can work backwards and learn that way. at this point im so confused, i need to know what the correct method is.

OpenStudy (anonymous):

your post vanished?

OpenStudy (loser66):

thanks for posting. got it

OpenStudy (anonymous):

uw!

OpenStudy (anonymous):

thanks, im working through it now. ill let you know if i have any questions.

OpenStudy (anonymous):

What's important here is the methodology on how to go about doing the mathematical induction. Pay attention to the "case of n=1" logic and the assumption of that it works for n=k. The reason that induction works is that k is 1 in the beginning, and then it works for "2". If it works for "2", it works for "3", and so on.

OpenStudy (anonymous):

If you use this general logic and the setting up of an induction proof and use this as a guide, you have a good chance of doing an induction problem on your own now.

OpenStudy (anonymous):

You show that it works for the case of n=1 and then you assume it works for n=k. Then you show that if you assume n=k and you can make it work for n=k+1, you've made your proof work. So, for n=1: [2(1)-1]^2 = 1 and (1(2(1)-1)(2(1)+1))/3 = 1 So it works for n=1. Now assume it works for n=k. Add [2(k+1)-1]^2 to both sides. For the right side: (k(2k-1)(2k+1))/3 + [2(k+1)-1]^2 = (k(2k-1)(2k+1))/3 + 3([2(k+1)-1]^2)/3 = (4k^3 - k)/3 + 3[(2k + 1)^2]/3 = (4k^3 - k)/3 + (12k^2 + 12k + 3)/3 = (4k^3 + 12k^2 + 11k + 3)/3 = (k + 1)(2k + 1)(2k + 3)/3 = ((k + 1)[2(k + 1) - 1][2(k + 1) + 1])/3 But this is just what we already assumed but now for the case n = k + 1, so we have now completed our proof.

OpenStudy (anonymous):

what about the left side?

OpenStudy (anonymous):

You added: [2(k+1)-1]^2 to both sides. You don't have to do anything to the left side. It's just the continuation of the "k" case but now for k + 1.

OpenStudy (anonymous):

It's just the next term in the sequence. It's the right side you prove with. The left side just shows the continuation.

OpenStudy (anonymous):

btw, Good luck to you in all of your studies and thx for the recognition! @comput313

OpenStudy (anonymous):

so the right side is 8k^2+2?

OpenStudy (anonymous):

You see, once you make the assumption that it works for n=k, you can use laws of math to add the same thing to both sides and thereby keep the equality. You just re-write the right side into a format, with steps, that show it's the n= k + 1 case.

OpenStudy (anonymous):

that comes out as 34=14??

OpenStudy (anonymous):

Is it starting to make sense now, @comput313 ? Or are you going through my steps at this time?

OpenStudy (anonymous):

i substituted in k=2 to test, and came out with 14=34.

OpenStudy (anonymous):

I substituted in k=2 and came up with 10 = 10. I just logged back on because Openstudy kept dropping my connection. Let's try to stay in touch on this problem until you are comfortable with it.

OpenStudy (anonymous):

For the left side i have (2k-1)^2+(2(k+1)-1)^2.